# Detecting Toxic Flow

### v1

@@ 2.3 Features to predict toxic trades
For a given client $c\in{\mathcal{A}}$, we use the following additional features: (i) cash $\mathfrak{C}^{c}_{t^{-}}$, (ii) inventory $\mathfrak{Q}^{c}_{t^{-}}$, (iii) volume of the order $q^{c}_{t}$, (iv) spread $\mathfrak{S}_{t^{-}}$ in the market just before the order arrives, (v) imbalance $\mathfrak{I}_{t^{-}}$ in the LOB just before the order arrives, (vi) transformed available volume in the best bid, (vii) transformed available volume in the best ask, (viii) last ask price at the time of trade, (ix) last bid price at the time of trade, (x) last midprice at the time of trade, (xi) total number of market updates since starting date, (xii) number of trades made by client $c$, (xiii) total number of trades made by all clients, (xiv) expanding-window volatility estimate of the mid-price, and (xv) proportion of previous sharp trades made by client $c$. These fifteen features account for both the state of the LOB, and the cash and inventory of the client. The remaining 168 features account for recent activity in the LOB. These are features (a)–(h) above measured for each of the seven intervals and for each of the three clocks to obtain a total of 8×7×3=1688$\times$ 7$\times$ 3=1688 × 7 × 3 = 168 features. Thus, for each client, we employ 15+3×8×7=18315+3$\times$ 8$\times$ 7=18315 + 3 × 8 × 7 = 183 features to predict the toxicity of their trades.
3 Algorithm for sequential estimation of toxicity
Let $({\mathcal{F}}_{t})_{t\geq 0}$ be the filtration that contains the information available to the broker. For a trade (and direction of trade) from client $c\in{\mathcal{A}}$ filled by the broker at time $t$, we denote by ${\bf x}_{t}\in\mathbb{R}^{M}$ the features observed at time $t$, where $M$ is the number of features, and the variable $y_{t}\in\{0,1\}$ denotes if a period $\mathfrak{G}$ after $t$ a trade is toxic ($y_{t}=1$), where $y_{t}$ is ${\mathcal{F}}_{t+\mathfrak{G}}$-measurable but it is not ${\mathcal{F}}_{t}$-measurable. We assume that the client buys or sells based on observed features ${\bf x}_{t}\in\mathbb{R}^{M}$ that are ${\mathcal{F}}_{t}$-measurable. For a given trading day we create the dataset of observations $(\mathcal{D}_{t_{i}})_{i\in I}$, where $\mathcal{D}_{t}=({\bf x}_{t},y_{t})$, $I=\{1,2,\dots,\,T\}$, and $T$ is the number of trades filled by the broker. For $i\in I$ we let $t_{i}$ be the time when the order reaches the broker. To simplify notation we use $\mathcal{D}_{i}$ instead of $\mathcal{D}_{t_{i}}$; thus, we refer to the dataset $(\mathcal{D}_{t_{i}})_{i\in I}$ as $(\mathcal{D}_{i})_{i\in I}$. For $n\in\mathbb{N}$ we let $\mathcal{D}_{1:n}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{n}\}$ be an ordered collection of observations indexed through time. Hence, for $t\leq T$ we have that $\mathcal{D}_{1:t}\subseteq\mathcal{D}_{1:T}$. We divide the dataset into a warmup dataset ${\cal D}_{\text{warmup}}$ and a deploy dataset ${\cal D}_{\text{deploy}}$. Here, ${\cal D}_{\text{warmup}}$ goes from 28 June 2022 to 29 July 2022, and ${\cal D}_{\text{deploy}}$ goes from 1 August 2022 to 21 October 2022.
We model the likelihood of a toxic trade given the features ${\bf x}_{t}$ as a Bernoulli random variable whose probability of a trade being toxic depends on the features ${\bf x}_{t}$ and a set of parameters ${\bm{\theta}}=({\bf w},{\boldsymbol{\psi}})$, that is
$p(y\,|\,\bm{\theta};{\bf x}_{t})=\text{Bern}\Big{(}y\,|\,\sigma({\bf w}^{\intercal}\,g({\bf x}_{t};{\boldsymbol{\psi}}))\Big{)},\qquad y\in\{0,1\}\,,$ (3.1)
[…]
$\text{Bern}(a\,|\,b)=b^{a}\,(1-b)^{1-a},\ a\in\{0,1\}\text{ and }b\in[0,1]\,.$
@@ 3 Algorithm for sequential estimation of toxicity
We refer to ${\bf w}\in\mathbb{R}^{L}$ as the last-layer parameters and we refer to ${\boldsymbol{\psi}}\in\mathbb{R}^{D}$ as the feature-transform parameters. The function $\sigma({\bf w}^{\intercal}\,g({\bf x}_{t};{\boldsymbol{\psi}}))$ is a NNet for classification where $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function. Figure 3 shows a graphical representation of PULSE when the NNet is a multilayered-perceptron (MLP).
[…]
@@
Our goal is to estimate the posterior distribution for ${\bm{\theta}}=({\bf w},{\boldsymbol{\psi}})$ in (3.1) after each new $y_{t}$ is observed, i.e., after observing if the trade is toxic. In practice, the dimension $D$ of the featured-transformed parameters and the dimension $M$ of the feature space satisfy $D\gg M$, and it is costly to update $\bm{\theta}$. Larsen et al., (2021) argue that one can train a linear projection of the weights of a neural network onto a linear subspace and obtain comparable performance to that when the full set of weights is used. Furthermore, (Murphy,, 2023, S. 17.3.5) discusses the importance of the last-layer parameters of a neural network to estimate accurately the posterior-predictive estimation of Bayesian neural networks (BNNs). Thus, to reap the benefits of last-layer methods and to have a model that can be deployed and adapted online with FX data we take two steps. One, model the posterior distribution of the last-layer parameters ${\bf w}$. Two, project the hidden-layer parameters ${\boldsymbol{\psi}}$ onto an affine subspace and model a posterior distribution of the parameters in the subspace. More precisely, we assume that ${\boldsymbol{\psi}}\in\mathbb{R}^{D}$ can be decomposed as an affine projection
@@
where ${\bf A}\in\mathbb{R}^{D\times d}$ is the fixed projection matrix, ${\bf z}\in\mathbb{R}^{d}$ are the projected weights such that $d\ll D$, and ${\bf b}\in\mathbb{R}^{D}$ is the offset term. With this projection, we rewrite (3.1) as $p(y\,|\,{\bf z},{\bf w};{\bf x}_{t})=\text{Bern}\Big{(}y\,|\,\sigma({\bf w}^{\intercal}\,g({\bf A}\,{\bf z}+{\bf b};\,{\bf x}_{t}))\Big{)}$, $y\in\{0,1\}$, and to simplify notation, we define $h({\bf z};{\bf x}_{t})=g({\bf A}\,{\bf z}+{\bf b};{\bf x}_{t})$ and write
@@
We estimate ${\bf A}$, ${\bf b}$, and assign prior distributions for ${\bf w}$ and $\bf z$ in the warmup stage, and we estimate the posterior distributions for ${\bf w}$ and ${\bf z}$ in the deploy stage. Figure 4 shows these two phases, which we explain in the following two subsections.
[…]
Figure 4: Warmup and deployment stages. We use all data available from times $t_{0}$ to $t_{\text{warmup}}$ to estimate $\bf A$ and $\bf b$. At $t_{\text{warmup}}$, we initialise the variational approximations $\phi_{0}({\bf w})$ and $\varphi_{0}({\bf z})$. Finally, for every $t>t_{\text{warmup}}$, we estimate ${\bf w}_{t}$ and ${\bf z}_{t}$.
3.1 Warmup phase: estimating the projection matrix and the offset term
Given the size of the dataset, we divide ${{\cal D}_{\text{warmup}}}$ into $B$ non-intersecting random batches $\mathcal{D}_{(1)},\ldots,\mathcal{D}_{(B)}$ such that
@@ 3.1 Warmup phase: estimating the projection matrix and the offset term
To estimate ${\bf b}$ and $\bf A$, we minimise the negative loss-function
$-\log p(\mathcal{D}\,|\,{\bm{\theta}})=-\sum_{n=1}^{N}\log p(y_{t}\,|\,{\bm{\theta}},{\bf x}_{t})$ (3.4)
via mini-batch stochastic gradient descent (SGD) over ${{\cal D}_{\text{warmup}}}$, where $\mathcal{D}$ is any random batch.
The vector ${\bf b}$ is given by
@@
The projection matrix ${\bf A}$ is found using singular-value decomposition (SVD) over the iterates of the SGD optimisation procedure. To avoid redundancy, we skip the first $n$ iterations and store the iterates every $k$ steps. The dimension of the subspace $d$ is found via hypeparameter tuning. Convergence to a local minimum of (3.4) is obtained through multiple passes of the data. Algorithm 1 shows the training procedure for a number $E$ of epochs.
1 def warmupParameters:
@@ 6    ${\bm{\theta}}^{(e)}\leftarrow{\bm{\theta}}^{(e-1)}-\alpha\,\bm{\kappa}(\text{gradient})$
In Algorithm 1, the parameter $\alpha$ is the learning rate, and the function $\bm{\kappa}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the per-step transformation of the Adam algorithm; see Kingma and Ba, (2015). At the end of the $E$ epochs, we obtain ${\bm{\theta}}^{(E)}=({\boldsymbol{\psi}}^{(E)},{\bf w}^{(E)})$. Then, the offset term ${\bf b}$ is given by
@@
3.2 Deploy phase: estimating last-layer parameters and feature-transform parameters
Given ${{\cal D}_{\text{deploy}}}$, we obtain a variational approximation to the posterior distributions of ${\bf w}$ and ${\bf z}$. We introduce Gaussian priors for both ${\bf w}$ and ${\bf z}$ at the beginning of the deploy stage. Next, let $t=0$ denote the last timestamp in the warmup dataset and $t=1$ the first timestamp of the deploy dataset, and let $\phi_{t}$ and $\varphi_{t}$ denote the posterior distribution estimates for ${\bf w}$ and ${\bf z}$ at time $t$. The initial estimates are given by
@@ 3.2 Deploy phase: estimating last-layer parameters and feature-transform parameters
For time $t\geq 1$, the variational posterior estimates are given by
${\phi_{t}\left({\bf w}\right)}$ $=\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}_{t}},\,{{\boldsymbol{\Sigma}}_{t}}\right)\quad\text{ and }\quad{\varphi_{t}\left({\bf z}\right)}=\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}_{t}},\,{{\boldsymbol{\Gamma}}_{t}}\right).$
Next, to find the posterior parameters ${\boldsymbol{\mu}}_{t},{\boldsymbol{\nu}}_{t},{\boldsymbol{\Gamma}}_{t},{\boldsymbol{\Sigma}}_{t}$, we recursively solve the following variational inference (VI) optimisation problem
${\boldsymbol{\mu}}_{t},\,{\boldsymbol{\nu}}_{t},\,{\boldsymbol{\Gamma}}_{t},\,{\boldsymbol{\Sigma}}_{t}=\operatorname*{arg\,min}_{{\boldsymbol{\mu}},\,{\boldsymbol{\nu}},\,{\boldsymbol{\Gamma}},\,{\boldsymbol{\Sigma}}}\text{KL}\left(\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)\,\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)||\phi_{t-1}({\bf w})\,\varphi_{t-1}({\bf z})\,p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})\right)\,,$ (3.6)
@@
for probability density functions $p$ and $q$ with the same support. The following theorem shows the update and prediction equations of the PULSE method.
@@
Suppose $\log p(y_{t}\,|\,{\bf z},{\bf w};{\bf x}_{t})$ is differentiable with respect to $({\bf z},{\bf w})$ and the observations $\{y_{t}\}_{t=1}^{T}$ are conditionally independent over $({\bf z},{\bf w})$. Write the mean of the target variable $y_{t}$ as a first-order approximation of the parameters centred around their previous estimate. Let $\sigma(x)=(1+\exp(-x))^{-1}$ be the sigmoid function and $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$ its derivative. Then, an approximate solution to (3.6) is given by
${\boldsymbol{\nu}}_{t}$ $={\boldsymbol{\nu}}_{t-1}+{\boldsymbol{\Sigma}}_{t-1}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\Big{(}y_{t}-\sigma({\boldsymbol{\nu}}_{t-1}^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t}))\Big{)},$ (3.7)
${\boldsymbol{\Sigma}}_{t}^{-1}$ $={\boldsymbol{\Sigma}}_{t-1}^{-1}+\sigma^{\prime}\big{(}{\boldsymbol{\nu}}_{t-1}^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\big{)}h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t}),$ (3.8)
${\boldsymbol{\mu}}_{t}$ $={\boldsymbol{\mu}}_{t-1}+{\boldsymbol{\Gamma}}_{t-1}\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\Big{(}y_{t}-\sigma({\boldsymbol{\nu}}_{t-1}^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t}))\Big{)},$ (3.9)
${\boldsymbol{\Gamma}}_{t}^{-1}$ $={\boldsymbol{\Gamma}}_{t-1}^{-1}+\sigma^{\prime}\big{(}{\boldsymbol{\nu}}_{t-1}^{\intercal}\,h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\big{)}\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{t-1};{\bf x}_{t})^{\intercal}\,,$ (3.10)

### v2

@@ 3.2 Features to predict toxic trades
For a given client $c\in{\mathcal{A}}$, we use the following additional features: (i) cash,(ii) inventory,(iii) volume of the order,(iv) spread in the market just before the order arrives, (v) imbalance in the LOB just before the order arrives, (vi) transformed available volume in the best bid, (vii) transformed available volume in the best ask, (viii) last ask price at the time of trade, (ix) last bid price at the time of trade, (x) last midprice at the time of trade, (xi) total number of market updates since starting date, (xii) number of trades made by client $c$, (xiii) total number of trades executed by all clients, (xiv) volatility estimate of the mid-price, and (xv) proportion of previous toxic trades executed by client $c$. We apply log-transformations to stabilise scale and limit outlier leverage: a signed log-transformation of the form $\text{sign}(x)\,\log(1+|x|)$ for (i) cash and (ii) inventory, and a $\log(1+x)$ transform for volumes in (iii), (vi), and (vii). These monotone transforms reduce heteroskedasticity and outlier influence and improve numerical conditioning for the downstream models; see e.g., West (2022). The fifteen features above account for both the state of the LOB, and the cash and inventory of the client. The remaining 168 features account for recent activity in the LOB. These are features (a)–(h) above, measured for each of the seven intervals and for each of the three clocks to obtain a total of $8\times 7\times 3=168$ features. Thus, for each client, we employ $15+8\times 7\times 3=183$ features to predict the toxicity of their trades.
In C.2 we illustrate that the performance of PULSE does not improve with additional information from the volume and transaction clock. In what follows we use the features provided by the three clocks to align with prior work and for completeness.
4 The PULSE method
For a trade from client $c\in{\mathcal{A}}$ filled at time $t$, let ${\bf x}_{t}\in\mathbb{R}^{M}$ denote the features observed at $t$ (with $M$ the number of features). Define $y_{t}\in\{0,1\}$ as the toxicity indicator evaluated $\mathcal{G}>0$ after $t$, with $y_{t}=1$ if the trade is toxic within the horizon, and 0 otherwise. Thus, ${\bf x}_{t}$ is observed at time $t$, whereas $y_{t}$ is known at time $t+\mathcal{G}$.
For each trading day we collect observations $(\mathcal{D}_{t_{i}})_{i\in I}$, where $\mathcal{D}_{t_{i}}=({\bf x}_{t_{i}},y_{t_{i}})$, $I=\{1,2,\dots,N\}$, $t_{i}$ is the arrival time of observation $i$, and $N$ is the number of trades that day. For notational convenience we write $\mathcal{D}_{i}$ for $\mathcal{D}_{t_{i}}$ and refer to the dataset as $(\mathcal{D}_{i})_{i\in I}$. For $n\in\mathbb{N}$, let $\mathcal{D}_{1:n}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{n}\}$ denote the first $n$ observations; in particular, for $n\leq N$ we have $\mathcal{D}_{1:n}\subseteq\mathcal{D}_{1:N}$.
Conditioned on the features ${\bf x}_{t}$, the toxicity indicator $y_{t}\in\{0,1\}$ is modelled as Bernoulli random variable with probability mass function
$p(y\,|\,\bm{\theta};{\bf x}_{t})=\text{Bern}\Big(y\,|\,\sigma({\bf w}^{\intercal}\,g({\boldsymbol{\psi}};{\bf x}_{t}))\Big),\qquad y\in\{0,1\}\,$, (4.1)
where $g:\mathbb{R}^{M}\to\mathbb{R}^{L}$ is the output-layer of a NNet, ${\bm{\theta}}=({\bf w},{\boldsymbol{\psi}})$ are the parameters of the neural network, $\bf w$} are the parameters of the last layer, and $\boldsymbol{\psi}$} are the parameters in the hidden layers. Here, and throughout the paper, we adopt the convention that $p$ denotes a likelihood function or a posterior density function. The function $\text{Bern}(\cdot\,|\,\cdot)$ is given by
$\text{Bern}(a\,|\,b)=b^{a}\,(1-b)^{1-a},\ a\in\{0,1\}\text{ and }b\in(0,1)\,.$
@@ 4 The PULSE method
We refer to ${\bf w}\in\mathbb{R}^{L}$ as the last-layer parameters and we refer to ${\boldsymbol{\psi}}\in\mathbb{R}^{D}$ as the feature-transform parameters. The function $\sigma({\bf w}^{\intercal}\,g({\boldsymbol{\psi}};{\bf x}_{t}))$ is a NNet for classification where $\sigma(x)=(1+\exp(-x))^{-1}$ is the sigmoid function. Figure 4 shows a graphical representation of the parameters that PULSE updates when the NNet is a multilayered-perceptron (MLP). Although we choose an MLP in the experiments, PULSE can be used with any NNet architecture.
$p(1|\sigma({\bf w}^{\intercal}g(\boldsymbol{{\bf x}_{t};\psi}))$${\bf x}_{t}\in\mathbb{R}^{M}\ldots\vdots\ldots\ldots$$\boldsymbol{\psi}$$\bf{w}$${\boldsymbol{\theta}}=({\bf w},{\boldsymbol{\psi}})$
@@
4.1 Sequential learning
We update the model parameters sequentially after each observed toxicity label to incorporate new information. Specifically, we update ${\bm{\theta}}=({\bf w},{\boldsymbol{\psi}})$ in (4.1) after each new $y_{t}$ is observed, i.e., after observing if the trade is toxic. In practice, the dimension $D$ of the featured-transformed parameters and the dimension $M$ of the feature space satisfy $D\gg M$, so it is costly to update $\bm{\theta}$ after each new observation using standard training techniques. Thus, the literature proposes various approaches to estimate the parameters of a NNet at a lower computational cost. In this paper, we build on two of these methods: lottery-ticket and last-layer methods.
Lottery-ticket methods exploit the over-parametrisation of NNets, in the sense that “a randomly-initialised dense NNet contains subnetworks that, when trained in isolation, reach test accuracy comparable to that of the original network”, Frankle and Carbin (2019). The lottery-ticket hypothesis states that such a subnetwork exists, and subnetworks satisfying the lottery-ticket hypothesis are called winning tickets. Winning tickets are linear projections of the NNet parameters onto a subspace; see Li et al. (2018) and Larsen et al. (2021). Duran-Martin et al. (2022) use the lottery-ticket hypothesis with a relatively small linear subspace, and use the extended Kalman filter (EKF) algorithm to propose a sequential update of the subspace parameters.
Alternatively, last-layer methods pre-train the NNet parameters in a warmup phase and then perform sequential updates on the last-layer parameters, see Murphy (2023, S. 17.3.5). Here, we employ both methods. Specifically, we propose a Bayesian approach to update the parameters of a NNet sequentially for both the subspace parameters in the hidden-layers and all of the parameters in the last layer. While previous literature focuses on updating either the last-layer or all parameters when performing online learning, ours is the first work that projects the parameters of the hidden layer and updates all of the units in the last layer. This decomposition enables full-rank updates of the last-layer parameters while restricting hidden-layer updates to a linear subspace. In doing so, it balances the rapid sequential updates typical of subspace neural networks with the statistical efficiency characteristic of last-layer methods.
In particular, we modify (4.1) and decompose the parameters of the hidden layer ${\boldsymbol{\psi}}\in\mathbb{R}^{D}$ as an affine projection of the form
@@ 4.1 Sequential learning
where ${\bf A}\in\mathbb{R}^{D\times d}$ is the fixed projection matrix and ${\bf z}\in\mathbb{R}^{d}$ are the projected (subspace) parameters such that $d\ll D$, and ${\bf b}\in\mathbb{R}^{D}$ is the offset term. The decomposition (4.2) provides a linear-subspace formulation of the lottery-ticket hypothesis and enables efficient updates of the hidden layer parameters. With this projection, we rewrite (4.1) as $p(y\,|\,{\bf z},{\bf w};{\bf x}_{t})=\text{Bern}\Big(y\,|\,\sigma({\bf w}^{\intercal}\,g({\bf A}\,{\bf z}+{\bf b};\,{\bf x}_{t}))\Big)$, $y\in\{0,1\}$. To simplify notation, we define $h({\bf z};{\bf x}_{t})=g({\bf A}\,{\bf z}+{\bf b};{\bf x}_{t})$ and write
@@
Our procedure has two stages: (i) an offline warmup phase that estimates $\bf A$} and $\bf b$} and uses ${\cal D}_{\text{warmup}}$ to select hyperparameters and (ii) an online phase that performs fast sequential updates of $({\bf w},{\bf z})$ on the live stream ${\cal D}_{\text{deploy}}$ while holding ${\bf A},{\bf b}$ fixed. Here, ${\cal D}_{\text{warmup}}$ is a one-off historical window, whereas ${\cal D}_{\text{deploy}}$ grows over time and is used for online learning and evaluation.
Figure 5 illustrates the workflow: an offline warmup phase using ${\cal D}_{\text{warmup}}$ (Section 4.1.1) followed by an online deploy phase using ${\cal D}_{\text{deploy}}$ (Section 4.1.2). The deploy phase proceeds indefinitely as trades arrive (the time index $T$ may grow over time).
$t_{0}$$t_{\text{warmup}}$TTInitialise $\phi_{0}({\bf w})$, $\varphi_{0}({\bf z})$warmup stageestimate $\bf A$, $\bf b$deploy stageestimate ${\phi_{t}\left({\bf w}\right)},{\varphi_{t}\left({\bf z}\right)}\,\;\forall t$
Figure 5: Warmup and deployment stages. We use all data available from $t_{0}$ to $t_{\text{warmup}}$ to estimate $\bf A$ and $\bf b$. At $t_{\text{warmup}}$, we initialise the variational approximations $\phi_{0}({\bf w})$ and $\varphi_{0}({\bf z})$. Finally, for $t>t_{\text{warmup}}$, we estimate ${\bf w}_{t}$ and ${\bf z}_{t}$.
4.1.1 Warmup phase: estimating the projection matrix and the offset term
This phase estimates $\bf A$} and $\bf b$}, and assigns prior distributions for $\bf w$} and $\bf z$. Given the size of the dataset, we divide ${\cal D}_{\text{warmup}}$} into $B$ non-intersecting random batches $\mathcal{D}_{(1)},\ldots,\mathcal{D}_{(B)}$ such that
@@ 4.1.1 Warmup phase: estimating the projection matrix and the offset term
To estimate $\bf b$} and $\bf A$, we use mini-batch stochastic gradient descent (SGD) over ${\cal D}_{\text{warmup}}$} to minimise the negative loss-function
$\log p(\mathcal{D}\,|\,{\bm{\theta}})=-\sum_{n=1}^{N}\log p(y_{n}\,|\,{\bm{\theta}},{\bf x}_{n})\,$, (4.4)
where $\mathcal{D}$ is any random batch. The vector $\bf b$} is given by
@@
Singular-value decomposition (SVD) over the iterates of the SGD optimisation procedure gives the projection matrix $\bf A$}. To avoid redundancy, we skip the first $n$ iterations and store the iterates every $k$ steps. The dimension of the subspace $d$ is found via hypeparameter tuning and convergence to a local minimum of (4.4) is obtained through multiple passes of the data. Algorithm 1 shows the training procedure for a number $E$ of epochs.
1def warmupParameters:
@@ 6    ${\bm{\theta}}^{(e)}\leftarrow{\bm{\theta}}^{(e-1)}-\alpha\,\bm{\kappa}(\text{gradient})$
In Algorithm 1, the parameter $\alpha$ is the learning rate, and the function $\bm{\kappa}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the per-step transformation of the Adam algorithm; see Kingma and Ba (2015). At the end of the $E$ epochs, we obtain ${\bm{\theta}}^{(E)}=({\boldsymbol{\psi}}^{(E)},{\bf w}^{(E)})$. Then, the offset term $\bf b$} is given by
@@
4.1.2 Deploy phase: online estimation of last-layer and subspace-feature-transform parameters
Here, we derive a novel, sample-free, and closed-form update rule that estimates the parameters of (4.3) sequentially. Specifically, in Proposition 1 in the Appendix, we find a set of fixed-point equations for the update rule. These equations need many iterations to converge and are computationally inefficient. Thus, Corollary 1 computes the gradient with respect to the subspace parameters to simplify the computations. Finally, Theorem 2 uses a Taylor expansion of the measurement model to obtain a closed-form solution to the set of fixed-point equations. This is computationally efficient because the update can be obtained in a single iteration.
We introduce Gaussian priors for both $\bf w$} and $\bf z$} at the beginning of the deploy stage. Let $n=0$ denote the last timestamp in the warmup dataset and $n=1$ the first timestamp of the deploy dataset. Denote by $\phi_{n}$ and $\varphi_{n}$ the posterior distribution estimates for $\bf w$} and $\bf z$} at time $t$, respectively. The initial estimates are given by
@@ 4.1.2 Deploy phase: online estimation of last-layer and subspace-feature-transform parameters
For $n\geq 1$, the variational posterior estimates are given by
${\phi_{n}\left({\bf w}\right)}$ $=\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}_{n}},\,{{\boldsymbol{\Sigma}}_{n}}\right)\quad\text{ and }\quad{\varphi_{n}\left({\bf z}\right)}=\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}_{n}},\,{{\boldsymbol{\Gamma}}_{n}}\right).$
Next, to find the posterior parameters ${\boldsymbol{\mu}}_{n},{\boldsymbol{\nu}}_{n},{\boldsymbol{\Gamma}}_{n},{\boldsymbol{\Sigma}}_{n}$, we recursively solve the following variational inference (VI) optimisation problem
${\boldsymbol{\mu}}_{n},\,{\boldsymbol{\nu}}_{n},\,{\boldsymbol{\Gamma}}_{n},\,{\boldsymbol{\Sigma}}_{n}=\operatorname*{arg\,min}_{{\boldsymbol{\mu}},\,{\boldsymbol{\nu}},\,{\boldsymbol{\Gamma}},\,{\boldsymbol{\Sigma}}}\text{KL}\left(\mathcal{N}\left({{\bf w}}\,|\,{{\boldsymbol{\nu}}},\,{{\boldsymbol{\Sigma}}}\right)\,\mathcal{N}\left({{\bf z}}\,|\,{{\boldsymbol{\mu}}},\,{{\boldsymbol{\Gamma}}}\right)||\phi_{n-1}({\bf w})\,\varphi_{n-1}({\bf z})\,p(y_{n}\,|\,{\bf z},{\bf w};{\bf x}_{n})\right)\,$, (4.6)
@@
for probability density functions $p$ and $q$ with the same support. The optimisation in (4.6) generalises the update rule for the Kalman filter when the parameters do not have a drift; see e.g., Lambert et al. (2021). The following theorem shows the update and prediction equations of the PULSE method.
@@
Suppose $\log p(y_{n}\,|\,{\bf z},{\bf w};{\bf x}_{n})$ is differentiable with respect to $({\bf z},{\bf w})$ and the observations $\{y_{n}\}_{n=1}^{N}$ are conditionally independent over $({\bf z},{\bf w})$. Write the mean of the target variable $y_{n}$ as a first-order approximation of the parameters centred around their previous estimate. Let $\sigma(x)=(1+\exp(-x))^{-1}$ be the sigmoid function and $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$ its derivative. Then, an approximate solution to (4.6) is given by
${\boldsymbol{\nu}}_{n}$ $={\boldsymbol{\nu}}_{n-1}+{\boldsymbol{\Sigma}}_{n-1}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})\Big(y_{n}-\sigma({\boldsymbol{\nu}}_{n-1}^{\intercal}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n}))\Big),$ (4.7)
${\boldsymbol{\Sigma}}_{n}^{-1}$ $={\boldsymbol{\Sigma}}_{n-1}^{-1}+\sigma^{\prime}\big({\boldsymbol{\nu}}_{n-1}^{\intercal}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})\big)h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})^{\intercal}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n}),$ (4.8)
${\boldsymbol{\mu}}_{t}$ $={\boldsymbol{\mu}}_{n-1}+{\boldsymbol{\Gamma}}_{n-1}\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})\Big(y_{n}-\sigma({\boldsymbol{\nu}}_{n-1}^{\intercal}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n}))\Big),$ (4.9)
${\boldsymbol{\Gamma}}_{n}^{-1}$ $={\boldsymbol{\Gamma}}_{n-1}^{-1}+\sigma^{\prime}\big({\boldsymbol{\nu}}_{n-1}^{\intercal}\,h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})\big)\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})\nabla_{{\bf z}}h({\boldsymbol{\mu}}_{n-1};{\bf x}_{n})^{\intercal}\,$, (4.10)